OpenStudy (hunter68):

The table below shows the movie preferences of high school and college students: Movie Preference Action Drama Comedy Row totals High School 40 23 12 75 College 20 51 28 99 Column totals 60 74 40 174 Julie mistakenly calculated the conditional relative frequency for college students who like comedy as being 70%. What did Julie actually calculate, and what is the correct answer?

OpenStudy (hunter68):

She calculated the joint relative frequency for college/comedy. The correct value of conditional relative frequency for college students who like comedy is 28%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct value of conditional relative frequency for college students who like comedy is16%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct answer is 28%. She calculated the joint relative frequency for college/comedy. The correct answer is 16%.

OpenStudy (hunter68):

@nikeboi101 @drawwithapurpose @Trevor12 @FROST_ZERO @nephycx

OpenStudy (hunter68):

please help...

OpenStudy (anonymous):

which one?

OpenStudy (hunter68):

what do you mean> which one i think it is?

OpenStudy (hunter68):

@-SpencerBraz- @Illuminated

OpenStudy (hunter68):

i think its C

OpenStudy (anonymous):

40 + 35 + 55 = 130 40/130 = .31 35/130 = .27 55/130 = .42

OpenStudy (hunter68):

im confused... is it C?

OpenStudy (anonymous):

yes

OpenStudy (hunter68):

okay thanks

OpenStudy (anonymous):

can i get a medal